# 1]organization=IIT-CNR, addressline=Via Moruzzi 1, code=56124, city=Pisa, country=Italy

### v3

@@ Abstract
Current content moderation follows a reactive, trial-and-error approach, where interventions are applied and their effects are only measured post-hoc. In contrast, we introduce a proactive, predictive approach that enables moderators to anticipate the impact of their actions before implementation. We propose and tackle the new task of predicting user abandonment following a moderation intervention. We study the reactions of 16,540 users to a massive ban of online communities on Reddit, training a set of binary classifiers to identify those users who would abandon the platform after the intervention—a problem of great practical relevance. We leverage a dataset of 13.8 million posts to compute a large and diverse set of 142 features, which convey information about the activity, toxicity, relations, and writing style of the users. We obtain promising results, with the best-performing model achieving micro F1-score $=0.914$. Our model shows robust generalizability when applied to users from previously unseen communities. Furthermore, we identify activity features as the most informative predictors, followed by relational and toxicity features, while writing style features exhibit limited utility. Theoretically, our results demonstrate the feasibility of adopting a predictive machine learning approach to estimate the effects of moderation interventions. Practically, this work marks a fundamental shift from reactive to predictive moderation, equipping platform administrators with intelligent tools to strategically plan interventions, minimize unintended consequences, and optimize user engagement.
@@ 1 Introduction
Online social media has greatly facilitated the sharing of information and opinions of all sorts, both locally and across the globe [1]. However, this also includes less desirable behavior, such as the spread of online toxic content, which besides potentially harming targeted users, is detrimental to the overall health and inclusivity of online environments [2, 3]. Online platforms are thus constrained to moderate such content, not only to maintain a safe and welcoming environment, but also because of legal compliance requirements [4]. Moderation strategies can be broadly divided into two main categories: hard and soft. The former consists of permanently removing users, groups, or posts that are violating community guidelines, such as posts deemed toxic. This type of intervention —also called deplatforming— can have unintended consequences because certain users perceive content removals as a form of censorship, especially if unmotivated. As a consequence, they may feel less prone to post again, or they could even decide to migrate to other more tolerant platforms tolerated [5, 6, 7]. This led to the proposal of the second broad form of moderation —soft moderation— which warns users about the harmfulness of a content without removing it [8, 9]. The growing consensus among the thriving content moderation literature is that gradual approaches that initially apply less punitive interventions, and that subsequently escalate the severity whenever required, are generally more effective than other solutions [10]. Nevertheless, hard interventions are still the most widely used form of content moderation. A well-known example of this kind is the so-called Great Ban, a massive deplatforming intervention enforced by Reddit in response to the rise of toxic and hateful speech on the platform, which involved the permanent ban of around 2,000 subreddits in June 2020 [11]. Among the banned subreddits was r/The_Donald, a community of Trump supporters that was one of the most popular political subreddits at the time [12]. Before the ban, the community had already faced two milder moderation interventions: a quarantine and a restriction that limited its visibility and the users that could serve as moderators of the community [13]. This sequence of interventions has been the subject of several studies that assessed changes in activity levels, use of language, political bias, factual news sharing, and toxicity [13, 14, 15, 16, 17]. These works found that many users were not much affected by the interventions, showing no significant change in behavior afterward. However, others seemed to be resentful of the platform and significantly decreased their activity, so much so that some completely abandoned the platform [11, 18].
@@
The previous results highlight the complexity of content moderation. First, in content moderation one size does not fit all, and effective moderation interventions should necessarily be tailored to the targets of the moderation [19, 20]. Furthermore, the current moderation strategy follows a trial-and-error approach, where platforms apply sequences of interventions until they meet the desired outcome, or force the misbehaving users out of the platform. r/The_Donald was a prime example of this strategy, as it was subject to two unsuccessful interventions prior to the Great Ban [13]. Given that the first of such interventions was applied in June 2019, while the last in June 2020, it took several months for the platform to evaluate the effects of each intervention and to impose further restrictions. Importantly, during that time the misbehaving users continued to spread toxic content and to engage in harmful behaviors.
@@ 1.1 Contributions
Motivated by the need for a strategic planning and deployment of moderation actions, we introduce and tackle the new task of predicting the effect of a moderation intervention. While many studies carried out post-hoc descriptive analyses of the effects of moderation interventions [15, 9, 16], here for the first time we adopt a predictive approach. We leverage the results of previous studies on the Great Ban as our ground-truth for the effects of that intervention [21, 11]. Then, we analyze user behavioral traces before the Great Ban to predict possible changes in the behavior of those users after the ban. We specifically focus on predicting those users who abandon the platform after the ban, in a binary classification task. This choice is both theoretically and practically relevant. Changes in user activity are among the most widely studied effects of past moderation interventions [13, 21, 22]. As such, from the theoretical standpoint it is interesting to assess the extent to which those effects are predictable. Additionally, the task is also practically relevant since the loss of users negatively affects platform revenues, as user engagement and popularity are paramount in the economic success of social media [23, 24]. Moreover, the abandonment of influential users can also trigger chain reactions, as their followers may also choose to leave the platform, further worsening the issue. The main results of this work are summarized in the following:
  • We proposed the new task of predicting users who abandon the platform following a moderation intervention and we evaluated different classification algorithms, hyper-parameters, and features sets for the task. Our best classifier is capable of detecting abandoning and non-abandoning users with a promising micro F1 score $=0.914$.
@@ 2.1 Descriptive moderation
In discussing the many works that took a descriptive approach to analyzing moderation interventions, we first survey studies on the Great Ban and other hard interventions, because they are the most similar to our present work. Among the works that studied the Great Ban is [21], which described the changes in activity and in-group vocabulary of those users who participated in the 15 most popular subreddits out of the 2,000 shut with the ban, using a word frequency divergence. The results highlighted the heterogeneity of the effects of the intervention and that top users tended to reduce their activity and their use of in-group vocabulary. The studies in [11, 18] carried out similar analyses but focused on evaluating changes in toxicity rather than activity. The authors found that 15.6% of the affected users abandoned Reddit after the ban. Among those who stayed there was a general reduction in toxic comments, but a small subset of users drastically increased their toxicity. Other works that evaluated the effects of hard content moderation include [25, 16]. The former measured changes occurred in Google Trends and Wikipedia, using the Google Knowledge Graph identifiers to apply time-series analysis on the digital attention traces. The latter studied the deplatforming of Twitter influencers and its effects on the activity of their supporters. Both works revealed that banning toxic influencers reduced the attention toward them and the number of posts citing them. These studies concluded that deplatforming may be a valid moderation strategy when used appropriately.
@@
Due to its prominence on Reddit and the multiple interventions that it faced, r/The_Donald was the focus of many works that investigated the effectiveness of online content moderation. For example, some works showed how the quarantine was able to reduce the visibility of the subreddit. At the same time however, it was unsuccessful at reducing racist and misogynistic language [22, 15]. Others examined the interventions in terms of activity, toxicity, factual reporting, and political bias both at community and user levels, using interrupted time series (ITS) regression analysis and Bayesian structural time series (BSTS) modeling [13, 14]. Community-level results found a general reduction in activity, a strong long-term increase in toxicity, a slight decrease in factual reporting, and no particular change in political ideology [13]. However, user-level reactions were more diversified and sometimes even extreme, especially for the most active users [14]. These results highlight that community-level effects are not always representative of the underlying user-level effects, which once again reaffirms the limitations of one-size-fits-all moderation [19, 20]. On the contrary, studying, and possibly even predicting user-level reactions to moderation interventions could be particularly beneficial to moderators.
@@ 2.2 Predictive moderation
Recent computational efforts in the study of content moderation increasingly adopt predictive rather than descriptive approaches. Among this body of work, the vast majority of studies tackled the task of predicting which pieces of content would be subject to moderation. For example, [26] predicted with high accuracy which YouTube videos would be later moderated using video, comment, and text-related features. The predictions were accurate even at posting-time, reducing the disappointment users face upon deletion after publication. Similarly, [27] proposed Lambretta, a learning to rank system developed to identify tweets likely to be removed from the platform because they convey false information. The system is intended to provide support to moderators in preventing the spread of fake news. Along the same line, CrossMod [28] automatically detects which Reddit posts are likely to be removed by the moderators, using cross-community learning via an ensemble of classifiers. The system is also capable of taking actions based on certain conditions set by the moderators themselves. The aforementioned works serve a twofold goal. On the one hand, they contribute to reverse engineer the content moderation actions taken by large online platforms, which often lack transparency and consistency [4]. On the other hand, the automated systems developed as part of the studies could also support human moderators by reducing the burden of extensive manual analyses as well as by reducing the emotional toll that comes from being exposed to toxic, harmful, or otherwise inappropriate content [29]. Other related works focused on predicting user behaviors, such as [30] that aimed to identify those users that would evade a Wikipedia ban by creating a new account. Additional tasks in the same work included early detection of such evasion and account matching. Instead, [31] discussed the feasibility of proactive moderation interventions on Reddit, looking at vocabulary features to identify problematic communities. Results showed that Reddit communities are constantly evolving, and that communities bound to become toxic can be preemptively identified. This early detection opens up the possibility to intervene before the problem escalates, thus possibly resulting in the application of less restrictive and punitive interventions.
@@ 2.3 Churn prediction
Churn prediction is a well-studied task in customer management that aims at forecasting whether a user will stop engaging with a platform or service based on their past behavior. It is related to our novel task of predicting user abandonment following a moderation intervention since the latter can be seen as a specific case of churn influenced by platform enforcement actions. Churn prediction has been traditionally tackled through various approaches, including the use of different feature sets, data sampling techniques, and machine learning models [32]. The majority of works tackled churn prediction in the context of telecom services. Some of them leverage Long Short-Term Memory (LSTM) and Convolutional Neural Networks (CNN) [33, 34]. The former does not consider information sequence by using a Bidirectional LSTM, while the latter uses temporal features with different granularity to maximize the accuracy. Other studies in this domain addressed data skewness through balancing techniques [35] or explored model explainability [36]. Other works study user adoption and churn related to the use of web browsers [37], by using a new model called Multivariate Behavior Sequence Transformer (MBST) to explore temporal and behavioral information separately. Another set of works covers churn prediction for the usage of financial services [38, 39], the former used CNNs to analyze text, while the second work used unstructured data like customer interactions as features. Although no prior research specifically addressed user retention in social media, numerous studies have examined users lifecycles within online communities. This body of work investigated various aspects of user engagement, such as how individuals interact with platforms [40] or how the way their language changes over time could be an indicator of the platform evolution [41]. Additionally, research has explored the motivations and patterns behind user migration to alternative platforms by analyzing user comments [42]. Temporal user dynamics on social media have been a key focus in studies such as [43, 44]. The former traced user trajectories from their initial joining phase as newcomers, to their eventual departure from a platform. Instead, the latter investigated the evolution of coordinated communities over time during a campaign by evaluating user retention in single communities. Lastly, Cunha et al. highlight how variations in both user and community behaviors contribute to different levels of community success [45]
@@ 3.1 Problem framing
$C_{+1}$ =$=\{u\;|\;A_{u}^{-}>0\;\text{and}\;A_{u}^{a}=0\}\quad\text{({abandoning} users)}$)
[…]
@@ 3.2 Task specification
$C^{H}$ :=$:=\{C\;|\;t_{+a}=t_{0}\;\implies\;P^{a}=P^{+}\}\quad\text{({hard abandonment})}$
$C^{S}$ :=$:=\{C\;|\;t_{+a}\gg t_{0}\;\implies\;P^{a}\subset P^{+}\}\quad\text{({soft abandonment})}.$
@@
Based on previous research that analyzed the medium-term effects of moderation interventions on Reddit [13, 11], we determined the duration of $P^{-}$ and $P^{+}$ to be 210 days, for a total time frame of 421 days (including the day $t_{0}$ of the Great Ban). The selected time frame provides a convenient and practical period that can be divided into months (7 months) or weeks (30 weeks). Hence, for $C^{H}$ the duration of $P^{a}=P^{+}=$ 7 months. On the other hand, for $C^{S}$ we selected the offset $t_{+a}=$ 4 months, with the duration of the corresponding abandonment period $P^{a}=$ 3 months (92 days), based on a precursory analysis of different offsets and class balances between non-abandoning and abandoning users. The 4-month inactivity threshold balances a long enough time to indicate meaningful disengagement while avoiding possible short-term fluctuations in user activity.

### v4

@@ Abstract
Current content moderation follows a reactive, trial-and-error approach, where interventions are applied and their effects are only measured post-hoc. In contrast, we introduce a proactive, predictive approach that enables moderators to anticipate the impact of their actions before implementation. We propose and tackle the new task of predicting user abandonment following a moderation intervention. We study the reactions of 16,540 users to a massive ban of online communities on Reddit, training a set of binary classifiers to identify those users who would abandon the platform after the intervention—a problem of great practical relevance. We leverage a dataset of 13.8 million posts to compute a large and diverse set of 142 features, which convey information about the activity, toxicity, relations, and writing style of the users. We obtain promising results, with the best-performing model achieving micro F1-score =0.914=0.914. Our model shows robust generalizability when applied to users from previously unseen communities. Furthermore, we identify activity features as the most informative predictors, followed by relational and toxicity features, while writing style features exhibit limited utility. Theoretically, our results demonstrate the feasibility of adopting a predictive machine learning approach to estimate the effects of moderation interventions. Practically, this work marks a fundamental shift from reactive to predictive moderation, equipping platform administrators with intelligent tools to strategically plan interventions, minimize unintended consequences, and optimize user engagement.
@@ 1 Introduction
Online social media has greatly facilitated the sharing of information and opinions of all sorts, both locally and across the globe [dwivedi2018social]. However, this also includes less desirable behavior, such as the spread of online toxic content, which besides potentially harming targeted users, is detrimental to the overall health and inclusivity of online environments [abarna2022identification, saumya2024filtering]. Online platforms are thus constrained to moderate such content, not only to maintain a safe and welcoming environment, but also because of legal compliance requirements [trujillo2023dsa, tessa2025improving, shahi2025year]. Moderation strategies can be broadly divided into two main categories: hard and soft. The former consists of permanently removing users, groups, or posts that are violating community guidelines, such as posts deemed toxic. This type of intervention —also called deplatforming— can have unintended consequences because certain users perceive content removals as a form of censorship, especially if unmotivated. As a consequence, they may feel less prone to post again, or they could even decide to migrate to other more tolerant platforms tolerated [gorwa2020algorithmic, jhaver2019does, horta2021platform]. This led to the proposal of the second broad form of moderation —soft moderation— which warns users about the harmfulness of a content without removing it [singhal2023sok, zannettou2021won]. The growing consensus among the thriving content moderation literature is that gradual approaches that initially apply less punitive interventions, and that subsequently escalate the severity whenever required, are generally more effective than other solutions [kiesler2012regulating]. Nevertheless, hard interventions are still the most widely used form of content moderation. A well-known example of this kind is the so-called Great Ban, a massive deplatforming intervention enforced by Reddit in response to the rise of toxic and hateful speech on the platform, which involved the permanent ban of around 2,000 subreddits in June 2020 [cima2024great]. Among the banned subreddits was r/The_Donald, a community of Trump supporters that was one of the most popular political subreddits at the time [massachs2020roots]. Before the ban, the community had already faced two milder moderation interventions: a quarantine and a restriction that limited its visibility and the users that could serve as moderators of the community [trujillo2022make]. This sequence of interventions has been the subject of several studies that assessed changes in activity levels, use of language, political bias, factual news sharing, and toxicity [trujillo2022make, trujillo2023one, shen2022tale, jhaver2021evaluating, etta2022comparing]. These works found that many users were not much affected by the interventions, showing no significant change in behavior afterward. However, others seemed to be resentful of the platform and significantly decreased their activity, so much so that some completely abandoned the platform [cima2024great, cima2025investigating].
@@
The previous results highlight the complexity of content moderation. First, in content moderation one size does not fit all, and effective moderation interventions should necessarily be tailored to the targets of the moderation [cresci2022personalized, cima2025contextualized]. Furthermore, the current moderation strategy follows a trial-and-error approach, where platforms apply sequences of interventions until they meet the desired outcome, or force the misbehaving users out of the platform. r/The_Donald was a prime example of this strategy, as it was subject to two unsuccessful interventions prior to the Great Ban [trujillo2022make]. Given that the first of such interventions was applied in June 2019, while the last in June 2020, it took several months for the platform to evaluate the effects of each intervention and to impose further restrictions. Importantly, during that time the misbehaving users continued to spread toxic content and to engage in harmful behaviors.
@@ 1.1 Contributions
Motivated by the need for a strategic planning and deployment of moderation actions, we introduce and tackle the new task of predicting the effect of a moderation intervention. While many studies carried out post-hoc descriptive analyses of the effects of moderation interventions [shen2022tale, zannettou2021won, jhaver2021evaluating], here for the first time we adopt a predictive approach. We leverage the results of previous studies on the Great Ban as our ground-truth for the effects of that intervention [trujillo2021echo, cima2024great]. Then, we analyze user behavioral traces before the Great Ban to predict possible changes in the behavior of those users after the ban. We specifically focus on predicting those users who abandon the platform after the ban, in a binary classification task. This choice is both theoretically and practically relevant. Changes in user activity are among the most widely studied effects of past moderation interventions [trujillo2022make, trujillo2021echo, chandrasekharan2022quarantined]. As such, from the theoretical standpoint it is interesting to assess the extent to which those effects are predictable. Additionally, the task is also practically relevant since the loss of users negatively affects platform revenues, as user engagement and popularity are paramount in the economic success of social media [carlson2020you, habib2022exploring]. Moreover, the abandonment of influential users can also trigger chain reactions, as their followers may also choose to leave the platform, further worsening the issue. The main results of this work are summarized in the following:
  • We proposed the new task of predicting users who abandon the platform following a moderation intervention and we evaluated different classification algorithms, hyper-parameters, and features sets for the task. Our best classifier is capable of detecting abandoning and non-abandoning users with a promising micro F1 score =0.914=0.914.
@@ 2.1 Descriptive moderation
In discussing the many works that took a descriptive approach to analyzing moderation interventions, we first survey studies on the Great Ban and other hard interventions, because they are the most similar to our present work. Among the works that studied the Great Ban is [trujillo2021echo], which described the changes in activity and in-group vocabulary of those users who participated in the 15 most popular subreddits out of the 2,000 shut with the ban, using a word frequency divergence. The results highlighted the heterogeneity of the effects of the intervention and that top users tended to reduce their activity and their use of in-group vocabulary. The studies in [cima2024great, cima2025investigating] carried out similar analyses but focused on evaluating changes in toxicity rather than activity. The authors found that 15.6% of the affected users abandoned Reddit after the ban. Among those who stayed there was a general reduction in toxic comments, but a small subset of users drastically increased their toxicity. Other works that evaluated the effects of hard content moderation include [ribeiro2024deplatforming, jhaver2021evaluating]. The former measured changes occurred in Google Trends and Wikipedia, using the Google Knowledge Graph identifiers to apply time-series analysis on the digital attention traces. The latter studied the deplatforming of Twitter influencers and its effects on the activity of their supporters. Both works revealed that banning toxic influencers reduced the attention toward them and the number of posts citing them. These studies concluded that deplatforming may be a valid moderation strategy when used appropriately.
@@
Due to its prominence on Reddit and the multiple interventions that it faced, r/The_Donald was the focus of many works that investigated the effectiveness of online content moderation. For example, some works showed how the quarantine was able to reduce the visibility of the subreddit. At the same time however, it was unsuccessful at reducing racist and misogynistic language [chandrasekharan2022quarantined, shen2022tale]. Others examined the interventions in terms of activity, toxicity, factual reporting, and political bias both at community and user levels, using interrupted time series (ITS) regression analysis and Bayesian structural time series (BSTS) modeling [trujillo2022make, trujillo2023one]. Community-level results found a general reduction in activity, a strong long-term increase in toxicity, a slight decrease in factual reporting, and no particular change in political ideology [trujillo2022make]. However, user-level reactions were more diversified and sometimes even extreme, especially for the most active users [trujillo2023one]. These results highlight that community-level effects are not always representative of the underlying user-level effects, which once again reaffirms the limitations of one-size-fits-all moderation [cresci2022personalized, cima2025contextualized]. On the contrary, studying, and possibly even predicting user-level reactions to moderation interventions could be particularly beneficial to moderators.
@@ 2.2 Predictive moderation
Recent computational efforts in the study of content moderation increasingly adopt predictive rather than descriptive approaches. Among this body of work, the vast majority of studies tackled the task of predicting which pieces of content would be subject to moderation. For example, [kurdi2020video] predicted with high accuracy which YouTube videos would be later moderated using video, comment, and text-related features. The predictions were accurate even at posting-time, reducing the disappointment users face upon deletion after publication. Similarly, [paudel2023lambretta] proposed Lambretta, a learning to rank system developed to identify tweets likely to be removed from the platform because they convey false information. The system is intended to provide support to moderators in preventing the spread of fake news. Along the same line, CrossMod [chandrasekharan2019crossmod] automatically detects which Reddit posts are likely to be removed by the moderators, using cross-community learning via an ensemble of classifiers. The system is also capable of taking actions based on certain conditions set by the moderators themselves. The aforementioned works serve a twofold goal. On the one hand, they contribute to reverse engineer the content moderation actions taken by large online platforms, which often lack transparency and consistency [trujillo2023dsa]. On the other hand, the automated systems developed as part of the studies could also support human moderators by reducing the burden of extensive manual analyses as well as by reducing the emotional toll that comes from being exposed to toxic, harmful, or otherwise inappropriate content [steiger2021psychological]. Other related works focused on predicting user behaviors, such as [niverthi2022characterizing] that aimed to identify those users that would evade a Wikipedia ban by creating a new account. Additional tasks in the same work included early detection of such evasion and account matching. Instead, [habib2022proactive] discussed the feasibility of proactive moderation interventions on Reddit, looking at vocabulary features to identify problematic communities. Results showed that Reddit communities are constantly evolving, and that communities bound to become toxic can be preemptively identified. This early detection opens up the possibility to intervene before the problem escalates, thus possibly resulting in the application of less restrictive and punitive interventions.
@@ 2.3 Churn prediction
Churn prediction is a well-studied task in customer management that aims at forecasting whether a user will stop engaging with a platform or service based on their past behavior. It is related to our novel task of predicting user abandonment following a moderation intervention since the latter can be seen as a specific case of churn influenced by platform enforcement actions. Churn prediction has been traditionally tackled through various approaches, including the use of different feature sets, data sampling techniques, and machine learning models [geiler2022survey]. The majority of works tackled churn prediction in the context of telecom services. Some of them leverage Long Short-Term Memory (LSTM) and Convolutional Neural Networks (CNN) [khattak2023customer, alboukaey2020dynamic]. The former does not consider information sequence by using a Bidirectional LSTM, while the latter uses temporal features with different granularity to maximize the accuracy. Other studies in this domain addressed data skewness through balancing techniques [sikri2024enhancing] or explored model explainability [poudel2024explaining]. Other works study user adoption and churn related to the use of web browsers [wu2022customer], by using a new model called Multivariate Behavior Sequence Transformer (MBST) to explore temporal and behavioral information separately. Another set of works covers churn prediction for the usage of financial services [de2020incorporating, vo2021leveraging], the former used CNNs to analyze text, while the second work used unstructured data like customer interactions as features. Although no prior research specifically addressed user retention in social media, numerous studies have examined users lifecycles within online communities. This body of work investigated various aspects of user engagement, such as how individuals interact with platforms [malinen2015understanding] or how the way their language changes over time could be an indicator of the platform evolution [danescu2013no]. Additionally, research has explored the motivations and patterns behind user migration to alternative platforms by analyzing user comments [newell2016user]. Temporal user dynamics on social media have been a key focus in studies such as [yao2021join, tardelli2024temporal]. The former traced user trajectories from their initial joining phase as newcomers, to their eventual departure from a platform. Instead, the latter investigated the evolution of coordinated communities over time during a campaign by evaluating user retention in single communities. Lastly, Cunha et al. highlight how variations in both user and community behaviors contribute to different levels of community success [cunha2019all].
@@ 3.1 Problem framing
$C_{+1}$ $=\{u\;|\;A_{u}^{-}>0\;\text{and}\;A_{u}^{a}=0\}\quad\text{({abandoning} users)}$
$C_{-1}$ $=\{u\;|\;A_{u}^{-}>0\;\text{and}\;A_{u}^{a}>0\}\quad\text{({non-abandoning} users)}.$
@@ 3.2 Task specification
$C^{H}$ $:=\{C\;|\;t_{+a}=t_{0}\;\implies\;P^{a}=P^{+}\}\quad\text{({hard abandonment})}$
$C^{S}$ $:=\{C\;|\;t_{+a}\gg t_{0}\;\implies\;P^{a}\subset P^{+}\}\quad\text{({soft abandonment})}.$
@@
Based on previous research that analyzed the medium-term effects of moderation interventions on Reddit [trujillo2022make, cima2024great], we determined the duration of $P^{-}$ and $P^{+}$ to be 210 days, for a total time frame of 421 days (including the day $t_{0}$ of the Great Ban). The selected time frame provides a convenient and practical period that can be divided into months (7 months) or weeks (30 weeks). Hence, for $C^{H}$ the duration of $P^{a}=P^{+}=$ 7 months. On the other hand, for $C^{S}$ we selected the offset $t_{+a}=$ 4 months, with the duration of the corresponding abandonment period $P^{a}=$ 3 months (92 days), based on a precursory analysis of different offsets and class balances between non-abandoning and abandoning users. The 4-month inactivity threshold balances a long enough time to indicate meaningful disengagement while avoiding possible short-term fluctuations in user activity.